The sum of the angles of every triangle in a plane is 180 degrees. The measure of the largest angle of the triangle is 10 degrees more than one angle and 20 degrees more than the other. What is the measure of the largest angle?
Answer by mananth(16949)

the largest angle be x deg
one angle = x-10
second angle = x-20
..
x+x-10+x-20=180
3x-30=180
3x=180+30
3x=210
x=70 deg
the other angle will be 70-10=60deg
third angle = 70-20=50deg